OpenStudy (anonymous):

prove that every root of the polynomial \[\sum_{k=0}^{n}2^{k(n-k)}x ^{k}\] is real. If you solve this do not give me the solution, just give me a very very small hint. I don't want to waste a good problem.

OpenStudy (anonymous):

This is what I have got so far: I tried a lot of different things, but the thing that I realized is that this sum is recursive(took me a few hours to realize that), so I defined my \[f(a) = 1 + a*x\] then I defined a series of numbers as \[g_{n}\] such that \[g _{0} = f(0) = 1\]\[g _{1} = f(1) = 1+ x\]\[g _{2} = f(f(0) + f(1)) = 1+ 2x + x ^{2}\]\[g _{3} = f(3f(0) + f(3f(0) + f(1))) = 1+ 4x + 4x ^{2} + x^3\]\[g _{4} = f(7 + f(15 + f(7 + f(1)))\]\[g_{5} = f(15 + f(63 + f(63 + f(15 + f(1)))))\] and so on, I think I have got a pattern, it is easy to see as in the last one the number 15 defines the coefficient of the x^4 term. It is easy to write this form, but I am stuck here.

OpenStudy (anonymous):

\[\sum_{k=0}^{n}2^{k(n-k)}x^k\] this is polynomial

OpenStudy (anonymous):

this is the polynomial*

OpenStudy (empty):

\(z=z^*\)

OpenStudy (mathmate):

Use Descartes rule of signs to rule out any positive root (real or not). Sturm's theorem can help you find the number of negative real roots. See, for example, section 2 of: http://web.cs.iastate.edu/~cs577/handouts/polyroots.pdf